OpenStudy (anonymous):

For what values of k does y=5+3e^(kx) satisfy the differential equation dy/dx= 10-2y I get k=1 but the book says -2... My steps were: find y' and plug it into the left side of the equation and then plug y into the right side. This gave me 3e^(kx) = 3ke^(kx)... Where did I go wrong?

OpenStudy (anonymous):

But using your method: Since \(y=5+3e^{kx}\), you have \(y'=3ke^{kx}\). Subbing into the equation, you have \[3ke^{kx}=10-2(5+3e^{kx})\\ 3ke^{kx}=-6e^{kx}\\ 3k=-6\]

OpenStudy (anonymous):

I haven't formally learned diff eqs yet, taking it in the fall so I'm just doing the intro stuff in my calc book. And thanks, that's twice I have failed at distributing a simple constant! Gah!

OpenStudy (anonymous):

You're welcome!